Entries for Sale of Fixed Asset

Equipment acquired on January 8 at a cost of $147,100, has an estimated useful life of 15 years, has an estimated residual value of $9,550, and is depreciated by the straight-line method.

a. What was the book value of the equipment at December 31 the end of the fourth year?
$

Feedback

Book value is the initial cost of the fixed asset minus the accumulated depreciation.

b. Assuming that the equipment was sold on April 1 of the fifth year for 101,787.

1. Journalize the entry to record depreciation for the three months until the sale date. Round your answers to the nerest whole dollar if required.

Depreciation Expense-Equipment
Accumulated Depreciation-Equipment

Feedback

2. Journalize the entry to record the sale of the equipment. If an amount box does not require an entry, leave it blank. Do not round intermediate calculations.

Cash
Accumulated Depreciation-Equipment
Loss on Sale of Equipment
Equipment

2.

Depreciation by Two Methods

A Kubota tractor acquired on January 8 at a cost of $216,000 has an estimated useful life of ten years. Assuming that it will have no residual value.

a. Determine the depreciation for each of the first two years by the straight-line method.

First Year Second Year
$ $

b. Determine the depreciation for each of the first two years by the double-declining-balance method. Do not round the double-declining balance rate. If required, round your final answer to the nearest dollar.

First Year Second Year
$ $
Answer #1

Question 1

Requiement a

Straight line Method

A

Cost

$     1,47,100.00

B

Residual Value

$           9,550.00

C=A - B

Depreciable base

$     1,37,550.00

D

Life [in years left ]

                           15

E=C/D

Annual SLM depreciation

$           9,170.00

Year

Book Value

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$    1,47,100.00

$    9,170.00

$   1,37,930.00

$       9,170.00

2

$    1,37,930.00

$    9,170.00

$   1,28,760.00

$    18,340.00

3

$    1,28,760.00

$    9,170.00

$   1,19,590.00

$    27,510.00

4

$    1,19,590.00

$    9,170.00

$   1,10,420.00

$    36,680.00

Value at the end of forth year = $110420.00

.

Requirement b

Year

Book Value

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$    1,47,100.00

$    9,170.00

$   1,37,930.00

$       9,170.00

2

$    1,37,930.00

$    9,170.00

$   1,28,760.00

$    18,340.00

3

$    1,28,760.00

$    9,170.00

$   1,19,590.00

$    27,510.00

4

$    1,19,590.00

$    9,170.00

$   1,10,420.00

$    36,680.00

5

$    1,10,420.00

$    2,292.50

$   1,08,127.50

$    38,972.50

Year 5 april 1

Depreciation expense-Equipment

$         2,293

Accumulated Depreciation - Equipment

$        2,293

Question 2

Sale of Eqiuipment

Straight line

Cost

$   1,47,100.00

Accumulated depteciation

$      38,971.00

Book value

$   1,08,129.00

Sales price

$   1,01,787.00

Book value

$   1,08,129.00

Gain /(loss)

$      (6,342.00)

General Journal

Debit

Credit

Cash

$ 1,01,787.00

Accumulated Depreciation-Equipment

$      38,971.00

Loss on sale of Equipment

$        6,342.00

Equipment

$    1,47,100.00

(Equipment sold)

.

Requirement 2a

Straight line Method

A

Cost

$ 2,16,000.00

B

Residual Value

$                     -  

C=A - B

Depreciable base

$ 2,16,000.00

D

Life [in years left ]

                       10

E=C/D

Annual SLM depreciation

$     21,600.00

Year

Book Value

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$    2,16,000.00

$     21,600.00

$   1,94,400.00

$    21,600.00

2

$    1,94,400.00

$     21,600.00

$   1,72,800.00

$    43,200.00

.

Requirement 2b

Double declining Method

A

Cost

$ 2,16,000.00

B

Residual Value

$                     -  

C=A - B

Depreciable base

$ 2,16,000.00

D

Life [in years]

10

E=C/D

Annual SLM depreciation

$     21,600.00

F=E/C

SLM Rate

10.00%

G=F x 2

DDB Rate

20.00%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$    2,16,000.00

20.00%

$      43,200.00

$ 1,72,800.00

$          43,200.00

2

$    1,72,800.00

20.00%

$      34,560.00

$ 1,38,240.00

$          77,760.00